OpenStudy (anonymous):

HELP! FAN & MEDAL The sun produces 3.9 x 10^33 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 3.25 x 10^3 seconds?

OpenStudy (anonymous):

it's scientific notation I think but idk if i divide or multiply the two numbers?

OpenStudy (amorfide):

let me make it simpler if you had 3 ergs of radiant energy for each second, and you wanted to know how much you had in 4 seconds, you would have 3 per second 4 times, so 3x4 would give you your answer does this help you?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

7..?

OpenStudy (anonymous):

\(3.9\times 3.25=12.675\)

OpenStudy (anonymous):

\(10^{33+3}=10^{46}\)

OpenStudy (anonymous):

so ur answer would be \(12.675\times 10^{46}\)

OpenStudy (anonymous):

and then simplify

OpenStudy (anonymous):

\(1.2675\times 10^{47}\)

OpenStudy (anonymous):

got it?

OpenStudy (amorfide):

uhm 33+3 is definitely not 46

OpenStudy (anonymous):

oops i meant \(10^{36}\) and \(10^{37}\) up there sorry abou tthat

OpenStudy (mathstudent55):